# Evaluating Conversational Recommender Systems via Large Language Models

### v3

@@ 2.3. Large Langugae Models as Evaluators for Text Quality
Text quality evaluation plays a critical role in advancing natural language generation (NLG) tasks. Human evaluation remains an fundamental approach for assessing the quality of text generated by machine learning models (Guzmán et al., 2015; Gillick and Liu, 2010). However, its high cost, subjectivity and limited reproducibility pose significant challenges to ensuring fair and consistent comparisons across different natural NLP models and algorithms (Gillick and Liu, 2010; Clark et al., 2021; Karpinska et al., 2021b). In contrast, automated evaluation metrics such as BLEU and ROUGE offer advantages in terms of low cost and high reproducibility. However, empirical studies have shown that these metrics often exhibit poor alignment with human preferences (Reiter, 2018).
@@
Some work leveraged existing LLMs as evaluators, experimenting with with zero-shot prompting and in-context learning (ICL) methods, and have explored evaluation paradigms such as Likert-scale scoring and pair-wise comparison, as well as different output formats: having LLMs directly emit numeric scores, preferred candidates, or leveraging implicit output probabilities.  Chiang and Lee (2023b) introduced a method wherein they present LLMs with the exact human evaluation instructions, sample texts, and rating questions, and parse the models’ generated responses into numeric quality scores as a reproducible surrogate for expert human assessment.  Lin and Chen (2023) proposed a framework that leverages LLMs to automatically fill evaluation forms and score open-domain conversational system responses across multiple dimensions (e.g., grammar, relevance, and appropriateness).  Chen et al. (2023b) investigated the effectiveness of ChatGPT as the evaluator for reference-free text quality evaluation by comparing explicit numeric score generation, implicit probability-based scoring, and pairwise comparison paradigms, finding that explicit score generation delivers the most reliable performance across diverse NLG tasks.  Fu et al. (2024) introduced a LLM-based method leveraging zero-shot prompting and ICL to perform multi-aspect evaluation of texts by interpreting conditional generation probabilities as quality scores.  Liu et al. (2023) leveraged GPT-4’s chain-of-thoughts to auto-generate detailed evaluation steps and employs a form-filling paradigm with probability-weighted scoring to assess NLG output quality.  Chan et al. (2023) proposed a multi-agent debate framework that orchestrates a team of LLMs to collaboratively discuss and score generated responses, thereby mimicking human evaluation processes for NLG tasks.
@@ 3. Preliminaries
Based on previous practices in CRS evaluation (Wang et al., 2023a, b; Zhu et al., 2024b, a), in chit-chat-based conversational recommendation, we formalize a single interaction session as a conversation log
@@
Taking Figure 2 as an example, the interaction between user and system can be formalized as a conversation log $L={D,R}$, and the dialogue component
@@
[…]
@@ 4. CoRE: Our Proposed Method
In this section, we introduce our proposed framework, Conversational Recommendation Evaluator (CoRE). It consists of two sequential components: (1) LLM-As-Evaluator. Employing LLMs to evaluate CRS across 12 distinct factors that influence user experience one by one. This evaluation is based on the carefully designed assessment criteria, allowing the LLM to generate both scores and detailed reasons for each factor. (2) Multi-Agent Debater. Obtaining an overall evaluation score for the CRS through a multi-agent debate framework, which discusses and synthesizes the evaluation scores of the above 12 factors into a unified score.
@@
4.1. LLM-As-Evaluator for 12 Factors
Figure 3. The first part of our proposed evaluation framework encompasses four dimensions, covering a total of 12 factors.
@@ 4.1. LLM-As-Evaluator for 12 Factors
Table 1. The concepts of the 12 factors from four evaluation dimensions.
@@
Table 2. The instruction template we provide to the LLMs for 12-factor evaluation.
@@
Prior research has indicated that a range of factors may influence how users assess their interactions with the system, encompassing both recommendation quality and dialogue effectiveness. These factors include: (perceived) accuracy of recommendation (Contreras et al., 2021; Manzoor et al., 2024; Jin et al., 2021; Zhang et al., 2024a), diversity (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024a) of recommended items, novelty of recommended items (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024a), explainability of recommendation (Sakai, 2023; Jin et al., 2021; Zhang et al., 2024a, c), user intent understanding and response capability (e.g., maintaining a coherent dialogue flow, recovering from errors) (Sakai, 2023; Langevin et al., [n. d.]; Jin et al., 2021; Zhang et al., 2024a), the quality of system-generated text (e.g., fluency, linguistic appropriateness and naturalness) (Jin et al., 2021; Jannach, 2022; Langevin et al., [n. d.]), system informativeness for decision-making (e.g., information sufficiency, information accuracy) (Sakai, 2023; Jannach, 2022), proactiveness of the system (Kraus et al., 2020; Deng et al., 2025; Christakopoulou et al., 2016).
Inspired by the previous studies, we heuristically summarize and list 12 factors that affect user satisfaction with conversational recommender systems, and categorize them into four dimensions based on the evaluation subjects involved in user-system interactions. Table 1 presents the concept of each factor.
@@
The example in Figure 3 shows how the 12 factors correspond to evaluation subjects in the user-CRS interaction. The blue dashed line surrounds the dialogue between the user and the CRS, indicating that the dialogue actions dimension focuses on evaluating the conversational actions and strategies between the user and the CRS. In the example, the user’s query “Any movies similar to The Day After Tomorrow” expresses an intent to receive a movie recommendation, and the system correctly adopts the strategy of providing a recommendation in response. The orange dashed line surrounds the system-generated text, which refers to the language expression used by the system when responding to the user. In the illustration, the system responds using appropriate language in a relatively natural and grammatically correct manner. The purple dashed line surrounds the recommended items, showcasing the three movies recommended by the system: The Matrix, The Shawshank Redemption, and Avatar. These items are diverse in genre, but they are not novel enough and do not align with the user’s request for movies similar to The Day After Tomorrow. The green dashed line surrounds both the system’s text and the recommended items, emphasizing the response content dimension, which evaluates the relevance between the system’s text and the recommended items. For instance, the system’s response text has little correlation with the recommended movie list, resulting in a low semantic relevance.
To obtain evaluation scores for the 12 factors, we leverage LLMs as evaluators by providing them with dialogue logs and detailed evaluation guidelines for each factor, instructing them to generate scores ranging from 0 to 4 along with the corresponding reasoning for each factor. Table 2 illustrates the structure of the instruction provided to the LLM evaluators. Table 3 presents the evaluation guideline when FACTOR is set to Coherence, covering FACTOR_DEFINITION, EVALUATION_STANDARD, and EVALUATION_STEP. For details on the evaluation guideline for each factor, refer to Appendix A. An example of CONVERSATION_LOG can be found in Figure 9.
@@ 4.2. Multi-Agent Debater for Overall Evaluation
Even when we have collected metric scores across the 12 factors to evaluate a CRS, synthesizing these into a comprehensive assessment of the system’s overall effectiveness remains a challenging task. This section seeks to answer the following question: how can we derive a unified measure of overall performance for a CRS based on the evaluations from the 12 factors?
@@
Specifically, we heuristically define four representative agent roles, each emphasizing a distinct set of factors relevant to their expertise: Common User, Domain Expert, Linguist, and HCI Expert. The Common User prioritizes the system’s performance in terms of Effectiveness, Coherence, and Recoverability. The Domain Expert focuses on Novelty, Diversity, and Groundness. The Linguist evaluates aspects such as Appropriateness, Naturalness, and Grammatical Correctness. Lastly, the HCI Expert examines Semantic Relevance, Explainability, and Proactiveness.
For this classification, we provide the following rationale. Ordinary users prioritize the system’s efficiency in meeting their needs, expecting relevant recommendations and coherent dialogues that align with their intentions, ensuring a smooth experience and enhancing satisfaction (Nielsen, 2000; Sundar et al., 2018; Zhou et al., 2015). Domain experts, particularly movie experts, expect the system to offer valuable, diverse insights and are more sensitive to the accuracy of the information due to their specialized knowledge. For linguists, their primary concern may be the linguistic quality of the system’s responses, particularly its appropriateness, naturalness, and grammatical correctness, which directly overlaps with the previously defined Language Expression dimension. HCI experts, on the other hand, focus on semantic relevance, explainability, and proactivity, as these factors are crucial for creating user-centered systems and fostering a positive user experience.
@@
Appendix D provides additional details about the multi-agent debate framework, including the role-playing instructions for each character. The content of CHAT_BUFFER can be found in Table 28.
@@
Table 4. The instruction template we provide to the LLMs for multi-agent debate.

### v4

@@ 2.3. Large Language Models as Evaluators for Text Quality
Text quality evaluation plays a critical role in advancing natural language generation (NLG) tasks. Human evaluation remains a fundamental approach for assessing the quality of text generated by machine learning models (Guzmán et al., 2015; Gillick and Liu, 2010). However, its high cost, subjectivity and limited reproducibility pose significant challenges to ensuring fair and consistent comparisons across different natural language processing (NLP) models and algorithms (Gillick and Liu, 2010; Clark et al., 2021; Karpinska et al., 2021b). In contrast, automated evaluation metrics such as BLEU and ROUGE offer advantages in terms of low cost and high reproducibility. However, empirical studies have shown that these metrics often exhibit poor alignment with human preferences (Reiter, 2018).
@@
Some work leveraged existing LLMs as evaluators, experimenting with zero-shot prompting and in-context learning (ICL) methods, and have explored evaluation paradigms such as Likert-scale scoring and pair-wise comparison, as well as different output formats: having LLMs directly emit numeric scores or preferred candidates, or leveraging implicit output probabilities.  Chiang and Lee (2023b) introduced a method wherein they present LLMs with the exact human evaluation instructions, sample texts, and rating questions, and parse the models’ generated responses into numeric quality scores as a reproducible surrogate for expert human assessment.  Lin and Chen (2023) proposed a framework that leverages LLMs to automatically fill evaluation forms and score open-domain conversational system responses across multiple dimensions (e.g., grammar, relevance, and appropriateness).  Chen et al. (2023b) investigated the effectiveness of ChatGPT as the evaluator for reference-free text quality evaluation by comparing explicit numeric score generation, implicit probability-based scoring, and pairwise comparison paradigms, finding that explicit score generation delivers the most reliable performance across diverse NLG tasks.  Fu et al. (2024) introduced a LLM-based method leveraging zero-shot prompting and ICL to perform multi-aspect evaluation of texts by interpreting conditional generation probabilities as quality scores.  Liu et al. (2023) leveraged GPT-4’s chain-of-thoughts to auto-generate detailed evaluation steps and employs a form-filling paradigm with probability-weighted scoring to assess NLG output quality.  Chan et al. (2023) proposed a multi-agent debate framework that orchestrates a team of LLMs to collaboratively discuss and score generated responses, thereby mimicking human evaluation processes for NLG tasks.
@@ 3. Preliminaries
Based on previous practices in CRS evaluation (Wang et al., 2023a, b; Zhu et al., 2024a, b), in chit-chat-based conversational recommendation, we formalize a single interaction session as a conversation log
@@
Taking Figure 2 as an example, the interaction between user and system can be formalized as a conversation log L=D,RL={D,R}, and the dialogue component
@@
$r^{1}_{s}$ $=\{\text{``28 Years Later (2025)'', ``Sinners (2025)''}\},$
$r^{2}_{s}$ $=\{\text{``Death of a Unicorn (2025)'', ``The Monkey (2025)''}\},$
$r^{3}_{s}$ $=\{\text{``Y2K (2024)''}\}.$
@@ 4. CoRE: Our Proposed Method
In this section, we introduce our proposed framework, Conversational Recommendation Evaluator (CoRE). It consists of two sequential components: (1) LLM-As-Evaluator. Employing LLMs to evaluate CRS across twelve distinct factors that influence user experience one by one. This evaluation is based on the carefully designed assessment criteria, allowing the LLM to generate both scores and detailed reasons for each factor. (2) Multi-Agent Debater. Obtaining an overall evaluation score for the CRS through a multi-agent debate framework, which discusses and synthesizes the evaluation scores of the above twelve factors into a unified score.
@@
4.1. LLM-As-Evaluator for Twelve Factors
Figure 3. The first part of our proposed evaluation framework encompasses four dimensions, covering a total of twelve factors.
Table 1. The concepts of the twelves factors from four evaluation dimensions.
@@ 4.1. LLM-As-Evaluator for Twelve Factors
Table 2. The instruction template provided to the LLMs for 12-factor evaluation.
@@
Prior research has indicated that a range of factors may influence how users assess their interactions with the system, encompassing both recommendation quality and dialogue effectiveness. These factors include: (perceived) accuracy of recommendation (Contreras et al., 2021; Manzoor et al., 2024; Jin et al., 2021; Zhang et al., 2024c), diversity (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024c) of recommended items, novelty of recommended items (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024c), explainability of recommendation (Sakai, 2023; Jin et al., 2021; Zhang et al., 2024c, b), user intent understanding and response capability (e.g., maintaining a coherent dialogue flow, recovering from errors) (Sakai, 2023; Langevin et al., 2021; Jin et al., 2021; Zhang et al., 2024c), the quality of system-generated text (e.g., fluency, linguistic appropriateness and naturalness) (Jin et al., 2021; Jannach, 2022; Langevin et al., 2021), system informativeness for decision-making (e.g., information sufficiency, information accuracy) (Sakai, 2023; Jannach, 2022), proactiveness of the system (Kraus et al., 2020; Deng et al., 2025; Christakopoulou et al., 2016).
Inspired by the previous studies, we heuristically summarize and list twelve factors that affect user satisfaction with conversational recommender systems, and categorize them into four dimensions based on the evaluation subjects involved in user-system interactions. Table 1 presents the concept of each factor.
@@
The example in Figure 3 shows how the twelve factors correspond to evaluation subjects in the user-CRS interaction. The blue dashed line surrounds the dialogue between the user and the CRS, indicating that the dialogue actions dimension focuses on evaluating the conversational actions and strategies between the user and the CRS. In the example, the user’s query “Any movies similar to The Day After Tomorrow” expresses an intent to receive a movie recommendation, and the system correctly adopts the strategy of providing a recommendation in response. The orange dashed line surrounds the system-generated text, which refers to the language expression used by the system when responding to the user. In the illustration, the system responds using appropriate language in a relatively natural and grammatically correct manner. The purple dashed line surrounds the recommended items, showcasing the three movies recommended by the system: The Matrix, The Shawshank Redemption, and Avatar. These items are diverse in genre, but they are not novel enough and do not align with the user’s request for movies similar to The Day After Tomorrow. The green dashed line surrounds both the system’s text and the recommended items, emphasizing the response content dimension, which evaluates the relevance between the system’s text and the recommended items. For instance, the system’s response text has little correlation with the recommended movie list, resulting in a low semantic relevance.
To obtain evaluation scores for the twelve factors, we leverage LLMs as evaluators by providing them with dialogue logs and detailed evaluation guidelines for each factor, instructing them to generate scores ranging from 0 to 4 along with the corresponding reasoning for each factor. Table 2 illustrates the structure of the instruction provided to the LLM evaluators. Table 3 presents the evaluation guideline when FACTOR is set to Coherence, covering FACTOR_DEFINITION, EVALUATION_STANDARD, and EVALUATION_STEP. For details on the evaluation guideline for each factor, refer to Appendix D. An example of CONVERSATION_LOG can be found in Figure 13.
@@ 4.2. Multi-Agent Debater for Overall Evaluation
Even when we have collected metric scores across the twelve factors to evaluate a CRS, synthesizing these into a comprehensive assessment of the system’s overall effectiveness remains a challenging task. This section seeks to answer the following question: how can we derive a unified measure of overall performance for a CRS based on the evaluations from the twelve factors?
@@
Specifically, we heuristically define four representative agent roles, each emphasizing a distinct set of factors relevant to their expertise: Common User, Domain Expert, Linguist, and HCI Expert. The Common User prioritizes the system’s performance in terms of Effectiveness, Coherence, and Recoverability. The Domain Expert focuses on Novelty, Diversity, and Groundness. The Linguist evaluates aspects such as Appropriateness, Naturalness, and Grammatical Correctness. Lastly, the HCI Expert examines Semantic Relevance, Explainability, and Proactiveness. For this classification, we provide the following rationale.
Ordinary users prioritize the system’s efficiency in meeting their needs, expecting relevant recommendations and coherent dialogues that align with their intentions, ensuring a smooth experience and enhancing satisfaction (Nielsen, 2000; Sundar et al., 2018; Zhou et al., 2015).
Domain experts, particularly movie experts, expect the system to offer valuable, diverse insights and are more sensitive to the accuracy of the information due to their specialized knowledge. The Justification for introducing the Domain Expert is as follows. A specific challenge in evaluating CRSs is verifying the Groundness (factual veracity) of the generated item descriptions and attribute details. Average users often lack the encyclopedic knowledge required to detect subtle factual errors or fabricated details about niche items. A domain expert (e.g., a film critic or bibliophile) acts as a necessary gatekeeper for content accuracy. Furthermore, beyond accuracy, domain experts are uniquely positioned to evaluate the Novelty and Diversity of recommendations. They can distinguish between generic, popularity-biased suggestions and high-quality, serendipitous discoveries from the long tail, ensuring the system provides value beyond what users already know.
The Linguist is crucial for assessing communicative competence. Since user trust in CRS is heavily dependent on the naturalness and pragmatic appropriateness of the generated dialogue, this role ensures the system adheres to linguistic norms that lay users perceive but cannot rigorously quantify. Unlike traditional recommender systems that rely on static graphical interfaces, CRSs interact via Natural Language Generation (NLG). The quality of linguistic delivery is not merely a stylistic preference but a functional prerequisite for user trust. Research in human–computer conversation indicates that systems should understand and adopt users’ language, presenting information in a natural and logically ordered manner to facilitate smooth interactions (Langevin et al., 2021). Moreover, pragmatic competence, that is, the ability to use language appropriately in social contexts, is as critical as grammatical correctness (Davies, 2007). A Linguist is necessary to evaluate nuances that lay users often overlook or cannot articulate, such as tonal consistency, and sociolinguistic appropriateness. If a system generates accurate recommendations but employs unnatural phrasing or incoherent discourse markers, it risks diminishing the user’s perception of system intelligence and trustworthiness, thereby undermining the acceptance of the recommendations.
HCI experts, on the other hand, focus on semantic relevance, explainability, and proactivity, as these factors are crucial for creating user-centered systems and fostering a positive user experience. From a Human-Computer Interaction perspective, a CRS functions as an interactive decision-support system rather than a passive information retrieval tool. The evaluation of such systems requires a specialized focus on interaction dynamics that mediate user decision-making processes. An HCI expert is essential to assess critical usability dimensions such as Explainability (how well the system justifies its choices to foster trust) and Proactiveness (the system’s ability to employ mixed-initiative strategies to guide users). While a common user evaluates the final utility of an item, an HCI expert critiques the interaction flow and cognitive load. They ensure that the system’s dialogue strategy effectively reduces the user’s decision effort and provides sufficient transparency, which are determinants of long-term user satisfaction and system adoption.
@@
Appendix H provides additional details about the multi-agent debate framework, including the role-playing instructions for each character. The content of CHAT_BUFFER can be found in Table 38.
Table 4. The instruction template provided to the LLMs for multi-agent debate.
@@
Table 5. An example of {FACTOR_RESULT}